OpenStudy (j2lie):

Find vertex and range of y=-x^2-4x-4? How do you solve it?

OpenStudy (j2lie):

@whpalmer4 @agent0smith

OpenStudy (agent0smith):

vertex x coordinate: x = -b/(2a) then plug that into y=-x^2-4x-4 to get the y coordinate

OpenStudy (j2lie):

is it -1? shouldn't it be the vertex for the t-chart?

OpenStudy (agent0smith):

b=-4 and a=-1

OpenStudy (j2lie):

I put the wrong thing down.